# Planet formation regulated by galactic-scale interstellar turbulence

### v1

@@ Abstract
Planets form from protoplanetary discs of dust and gas that surround stars younger than a few million years.[1] The properties of these discs dictate how planets grow, determining the nature, architecture and composition of the observed exoplanet population. The nature of (turbulent) angular momentum transport through the disc in particular is critical for planet growth[2] and migration.[3, 4] While the dominant physical processes in disc evolution remain unknown, theoretical studies widely assume an isolated star-disc system. Here we challenge this conventional assumption by showing that, far from being isolated systems, discs in typical star forming environments are constantly replenished by capturing new material from the interstellar medium (ISM). We show that this in-fall alone explains observed disc masses, outer radii and stellar accretion rates as a function of time and stellar mass. ISM capture is also capable of driving disc turbulence corresponding to viscosity in the range $\alpha_{\mathrm{SS}}\sim 10^{-5}{-}10^{-1}$, as observationally inferred.[5, 6, 7] We find that $20{-}70$ percent of discs are composed of material captured in the most recent half of their life-time, implying their properties are not direct probes of internal disc physics. Our results suggest that planet formation is driven by the turbulent cascade of energy from galactic-scales down to the protoplanetary disc and that recent evidence of in-fall from the ISM onto mature protoplanetary discs[8, 9, 10] are part of an important, general process. This represents a far-reaching shift in our understanding of every stage of planet formation, from dust aggregation to the migration of giant planets through the disc.
ABSTRACT
MAIN
@@ ABSTRACT
Conventional models for planet formation assume that a protoplanetary disc forms during protostellar collapse, and then planets grow from material in the isolated star-disc system. This picture is now being challenged on a variety of fronts.[11, 12] Here we particularly explore the implications of recent Atacama Large Millimeter/submillimeter Array (ALMA) observations in molecular line emission[13] and in infrared scattered light[14, 15] that are frequently revealing extended structures associated with protoplanetary discs, on spatial scales several times the disc outer radius. In-fall has been suggested as a possible origin for these structures,[16, 17, 18] which are highly complex, including large scale spirals and streamers around T Tauri[19] and Herbig AeBe[20] stars, often associated with reflection nebulosity,[21] accretion outbursts[22, 23, 24] and misaligned inner discs.[25] Evidence that is suggestive of in-fall events has been found not only during the earliest stages of disc evolution, when the star is still embedded in its natal environment,[26, 27, 28] but also around stars such as AB Aur,[29, 30, 8] DR Tau[31] and SU Aur,[25] which are all mature star-disc systems of age $\sim 1{-}3$ Myr. A recent estimate from a sample size of $43$ disc-hosting stars suggests approximately $16$ percent are interacting with the ambient ISM.[32] This material should fall onto the disc on a time-scale much shorter than the disc lifetime,[16, 17] suggesting that either these events occur much more frequently than observed or the external structure is replenished by the ISM for a prolongued period. In-falling gas may be the consequence of Bondi-Hoyle-Lyttleton (BHL) accretion, where low velocity gas is captured due to the gravitational potential of the star.[33, 34] Theoretically, the rate of capture of material from the interstellar medium (ISM) via BHL accretion can be substantial while stars remain bound to a dense gas reservoir, and has been proposed to explain observed stellar accretion rates.[35, 36, 37] However, assessing the importance of this process for the observed population of protoplanetary discs throughout their lifetime requires quantifying the temporal evolution of the turbulent ISM surrounding forming stars.
In this work, we statistically quantify the time evolution of the BHL accretion rate. We summarise the approach here and refer to the Methods for details. First, we apply an excursion set formalism[38, 39] to model fluctuations in the gas density and velocity in the solar neighbourhood. We use this to draw a representative sample of gravitationally unstable star forming regions that are common within $\sim 200$ pc of the sun (Extended Data Figure 1). We compute the evolution of the radius, density and star formation rate in these regions as they undergo free-fall collapse.[40] Star formation continues until the fraction of gas converted into stars is $\epsilon=0.5$,[41] at which time the bound gas is dispersed. Having computed the evolution of a representative sample of star forming regions, we draw formation times for a representative sample of $500$ stars from these regions, weighted by the instantaneous star formation rate. We also draw individual star masses from a log-uniform distribution between $10^{-2}\,M_{\odot}$ and $3M_{\odot}$. Each star evolves within the potential of the collapsing star forming region from the time of formation until the dispersal of the region, at which point the density and velocity structure local to the star-disc system is determined by the evolution of the unbound ISM. Density and velocity fluctuations in the galactic ISM evolve at different scales according to a generalised Fokker-Planck equation, resulting in drift-diffusion behaviour. We model this via a random-walk,[39] and can therefore compute the BHL accretion rate, being defined at the scale which maximises the rate of capture of material onto the star-disc system. In this way, we construct accretion histories for all of our stars.
We illustrate the outcome of these calculations in Figure 1, for a subset of six stars in our sample. The phase in which the star remains within the bound gaseous reservoir typically persists for $\sim 1$ Myr. During this phase, the relative velocity of the gas is low, $\Delta v_{\mathrm{gas}}\lesssim 1$ km s-1 (Figure 1, top), in line with previous estimates.[37] Evidently, since the gas reservoir is gravitationally unstable, the local density $\rho_{\mathrm{gas}}$ (Figure 1, middle) is also large compared to the median density at the galactic mid-plane, which is $\rho_{\mathrm{mid}}\sim 10^{-24}$ g cm-3. This results in an early phase of enhanced BHL accretion at a median rate $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-8}\,M_{\odot}$ yr-1 (Figure 1, bottom). This median rate then quickly drops to $\dot{M}_{\mathrm{BHL,1/2}}\sim 10^{-10}{-}10^{-11}\,M_{\odot}$ yr-1 in the age range $1{-}3$ Myr for a solar mass star. This is considerably lower than the typical rate of accretion onto the central star, $\dot{M}_{\mathrm{acc}}\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1.[42] One might therefore assume that BHL accretion no longer plays an important role for disc evolution. However, in a turbulent medium, vorticity of the in-falling material means that accretion is mediated via a (protoplanetary) disc.[43, 17] In this case, the change in accretion rate can be averaged over the time-scale $\tau_{\mathrm{acc}}$ at which angular momentum is transported within the disc. Empirically, this time-scale can be a considerable fraction of the disc lifetime.[44] In this case the median underestimates the true contribution of BHL accretion rate onto the disc to the accretion rate onto the star. By contrast, the mean accretion rate when averaged over all stars in our model is $\langle\dot{M}_{\mathrm{BHL}}\rangle\sim 10^{-8}-10^{-9}\,M_{\odot}$ yr-1 at $1-3$ Myr, as shown as the dashed red line in Figure 1. However, the mean overestimates the typical BHL accretion rate contribution for individual discs, since it is dominated by a small number of rapid accretors. Interpreting the contribution from BHL accretion to the stellar accretion rate for typical star-disc systems is therefore not a simple task. Compounding this problem further, discs that survive to comprise the observed sample may also preferentially be those which have been replenished by the ISM. Interpreting the role of BHL accretion for disc evolution therefore requires further calculation.
To quantify the degree to which BHL accretion from the ISM sculpts observed protoplanetary disc properties, we perform a simple numerical experiment. For now, we remain agnostic to the mechanism by which accretion from the disc onto the star proceeds, parameterising the process by a single parameter $\tau_{\mathrm{acc}}=M_{\mathrm{disc}}/\dot{M}_{\mathrm{acc}}$, where $M_{\mathrm{disc}}$ is the instantaneous disc mass. Then the rate of change of the disc mass as a function of time $t$ is simply
$\dot{M}_{\mathrm{disc}}(t)=\dot{M}_{\mathrm{BHL}}(t)-\frac{M_{\mathrm{disc}}(t)}{\tau_{\mathrm{acc}}}.$
We assume that the initial disc mass is negligible, adopting $M_{\mathrm{disc}}(t=0)=0$ (see Methods). We then make a parametric choice, anchoring $\tau_{\mathrm{acc}}$ to observational constraints.[44] We draw a fixed $\tau_{\mathrm{acc}}$ for each star, adopting a log-uniform distribution with ±$\pm 1$ dex around $\log\tau_{\mathrm{acc}}/\mathrm{1\,Myr}=-0.5$. As a dispersal criterion, when $M_{\mathrm{disc}}<3\times 10^{-5}\,M_{\odot}$ then we assume the star would be considered disc-less in observations. Our results are not dependent on this choice within a sensible range, and we are motivated by the fact that few discs are detected below this threshold. This threshold also results in a fraction of surviving discs in agreement with observational constraints (Extended Data Figure 2). We further find that our calculations exhibit excellent agreement with the observed distributions of disc masses, stellar accretion rates, outer disc radii and constraints on turbulent velocities, as discussed below.
We show the disc properties in our model as a function of the stellar mass and age in Figure 2, comparing to observed discs in local star forming regions across a range of different ages. In order of increasing estimated age,[42] these regions are $\rho$ Ophiuchus, Taurus, Lupus, Chameleon I and Upper Scorpius (with the caveat that age estimates are uncertain and ages in Chameleon I, Taurus and Lupus overlap – see Methods). For nearly all these comparisons, we find remarkable agreement with the observed disc masses (Figure 2, top) and stellar accretion rates (Figure 2, middle) as a function of stellar host mass $m_{*}$ and age. Both are typically limited by an upper ‘envelope’ that is a super-linear function of stellar mass, approximately $\propto m_{*}^{2}$, with a broad distribution below this upper limit for any given $m_{*}$. For older stars ($\gtrsim 1{-}2$ Myr), the masses and accretion rates decrease as discs are no longer rapidly replenished by BHL accretion. Observed accretion rates for Taurus and Upper Scorpius deviate somewhat from our model outcome, and these moderate differences must originate from some variation in $\tau_{\mathrm{acc}}$ that we do not account for, since we choose to fix the $\tau_{\mathrm{acc}}$ distribution. There may be several reasons for such variation (see Methods), and the overall agreement of our simple model with observations is striking. Interestingly, the discs in our model typically take a few $10^{5}$ years to grow to a maximal mass (Extended Data Figure 3), which may explain the finding that the observationally inferred disc mass is not a monotonic function of age for very young star forming regions.[45, 46, 47]
We can also attempt a comparison with the disc outer radii in our model and from observations. While we do not directly compute the disc radius $R_{\mathrm{disc}}$ in our model, we do track the radius from which material accretes onto the disc $R_{\mathrm{acc}}$ (see Methods). We generally expect $R_{\mathrm{acc}}\gtrsim R_{\mathrm{disc}}$, since the resultant disc radius depends on the vorticity of the in-falling material during the BHL accretion process.[43] However, it is reasonable to expect that $R_{\mathrm{disc}}\propto R_{\mathrm{acc}}\propto m_{*}$, assuming that the vorticity does not depend on $m_{*}$ (or $R_{\mathrm{acc}}$). While there are fewer measurements of gaseous disc outer radii in local star forming regions, in the bottom row of Figure 2 we compare our model to the empirically-inferred relationship for the outer gas disc radius as inferred from CO molecular line measurements: $R_{\mathrm{CO}}=250\left(m_{*}/1\,M_{\odot}\right)^{0.9}$ au,[48] typically a factor $\sim 2{-}3$ larger than the dust component, although with considerable scatter and some examples of much more extended CO radii.[49] We again find a remarkable similarity between our model results and the observed disc properties, with a tail of large $R_{\mathrm{acc}}>R_{\mathrm{CO}}$.
The striking agreement between the model results and the observational disc properties strongly supports the conclusion that the outer gaseous disc radii, total disc masses and stellar accretion rates are largely set by accretion from the ISM. This finding is particularly compelling given the simplicity of the model, relying only on established physical processes, without requiring fine tuning of free parameters. However, what is less clear is whether these properties are set by the initial phase of rapid accretion, or whether disc material is constantly replenished throughout the disc lifetime. To explore this, we define $f_{\mathrm{M,1/2}}$, which is the fraction of the current disc mass that was accreted from the ISM within the most recent half of the disc life-time (see Methods). This fraction is shown as a function of time for our $500$ discs in Figure 3. We find that the distribution of $f_{\mathrm{M,1/2}}$ varies with time, with a greater fraction of discs being composed of recently accreted material early during their evolution. The fraction of discs that are not composed of any recently accreted material increases from approximately $20$ percent at $1$ Myr, to $50$ percent for discs older than $3$ Myr. This decrease is instigated by the drop in typical accretion rates at around $1$ Myr (Figure 1), after which the majority of stars no longer inhabit the collapsing molecular cloud in which they were born. The fraction of discs that are mostly composed of recently accreted material ($f_{\mathrm{M,1/2}}>0.5$) remains approximately in the range $20-70$ percent throughout the disc life-time. These estimates are conservative, since aside from stellar accretion we have not considered any internal processes, such as planet formation, that deplete disc material. We conclude that observed disc populations are constantly being replenished by the surrounding ISM, and that their properties therefore may not be direct probes of internal disc physics.
We have so far assumed that accretion onto the host star from the disc is driven by some, undefined internal angular momentum transport mechanism. However, we also expect some turbulence to be sustained due to the continuous in-fall of material. Material falls onto the disc at approximately the free-fall velocity, and then shocks as it hits the disc surface.[50] Some fraction $\epsilon_{\mathrm{t}}$ of the energy deposited into the disc is kinetic, and the resultant turbulence may be sufficient to drive the turbulent velocity $v_{\mathrm{t,disc}}$ empirically inferred for protoplanetary discs. We explore this possibility by computing the rate at which the turbulent energy evolves for our disc sample (see Methods). Turbulent energy decays rapidly within the disc, so BHL accretion onto the disc must be sustained. We can parameterise the strength of turbulence in terms of a Shakura-Sunyaev constant $\alpha_{\mathrm{SS}}$,[51] where $\alpha_{\mathrm{SS}}\sim v_{\mathrm{t,disc}}^{2}/c_{\mathrm{s}}^{2}$,[52] for sound speed $c_{\mathrm{s}}=0.2$ km s-1.
Assuming efficiency $\epsilon_{\mathrm{t}}=1$, we show the resulting evolution of the $v_{\mathrm{t,disc}}$ and $\alpha_{\mathrm{SS}}$ distribution in Figure 4. At early times $\lesssim 1$ Myr, our calculation would suggest that the turbulent velocity can actually exceed $c_{\mathrm{s}}$. Physically, this would produce shocks in addition to the initial shock from the free-fall onto the disc, and the resultant shock-heating would presumably substantially reduce $\epsilon_{\mathrm{t}}$. However, since the turbulent decay time-scale is short ($\sim 10^{4}$ years), $v_{\mathrm{t,disc}}$ remains in an almost instantaneously determined balance between in-fall and turbulent decay. Therefore any deviation from $v_{\mathrm{t,disc}}/c_{\mathrm{s}}$ for the early disc in our model do not influence the subsequent distribution. We can therefore compare the turbulent velocity distribution to observational constraints at any time without worrying about a complex dependence on a time-varying $\epsilon_{\mathrm{t}}$. We note also the weak dependence of $v_{\mathrm{t,disc}}$ on $\epsilon_{\mathrm{t}}$: $v_{\mathrm{t,disc}}\propto\epsilon_{\mathrm{t}}^{1/3}$, or $\alpha_{\mathrm{SS}}\propto\epsilon_{\mathrm{t}}^{2/3}$.
Observational constraints for the turbulent velocity in discs have yielded upper limits $v_{\mathrm{t,disc}}/c_{\mathrm{s}}\lesssim 0.05$ for older discs (age $\sim 5-10$ Myr) TW Hya,[53, 54] HD163296[55, 56], MWC 480[57, 6] and V4046 Sgr,[58, 6] and high turbulence for the younger discs around DM Tau[6] ($\sim 1{-}5$ Myr[57, 59]) and IM Lup[60] ($\sim 1$ Myr[61]) with $v_{\mathrm{t,disc}}/c_{\mathrm{s}}\sim 0.25-0.33$ and $v_{\mathrm{t,disc}}/c_{\mathrm{s}}\sim 0.4-0.6$ respectively, corresponding to an high $\alpha\sim 0.1$. One of the most stringent constraints on turbulence in the mid-plane of a disc has been inferred from the vertical scale of dust in edge-on discs. One example is Oph163131, for which the extremely well-settled dust suggests a very low $\alpha_{\mathrm{SS}}\sim 10^{-5}$.[62] Consistent with the constraints on the turbulent velocity from line broadening, young discs have been shown to exhibit much less vertical settling of mm-sized dust grains.[63, 64] The distribution of $\alpha_{\mathrm{SS}}$ that we obtain from our model imply that both discs with $\alpha_{\mathrm{SS}}\sim 0.1$ (similar to DM Tau) and at the other extreme $\alpha_{\mathrm{SS}}\sim 10^{-5}$ (similar to Oph163131) are possible if $\epsilon_{\mathrm{t}}$ is just somewhat below unity. Our model is also broadly consistent with the hint that older discs have lower levels of turbulence. We conclude that accretion from the ISM is capable of driving protoplanetary disc turbulence at the levels inferred for individual discs, providing $\epsilon_{\mathrm{t}}\sim 0.1{-}1$.
A related question is whether the turbulence injected from the ISM is sufficient to drive observed accretion rates. To avoid adding prohibitive complexity to our model (see discussion in Methods), in this work we fix the accretion time-scale $\tau_{\mathrm{acc}}$ (or turbulence) for each star-disc system, such that it is independent of the environment. However, we can estimate whether the capture of ISM material could feasibly drive turbulence at the same level that we assume in our distribution of $\tau_{\mathrm{acc}}$. To do so, we first rewrite[65]
$\tau_{\mathrm{acc}}=\frac{M_{\mathrm{disc}}}{\dot{M}_{\mathrm{acc}}}=\frac{2R_{1}^{2}\Omega_{1}}{3\alpha_{\mathrm{SS}}c_{\mathrm{s}}^{2}}=0.37\,\left(\frac{\alpha_{\mathrm{SS}}}{0.1}\right)^{-1}\left(\frac{R_{1}}{250\,\mathrm{au}}\right)^{1/2}\,\left(\frac{m_{*}}{1\,M_{\odot}}\right)^{1/2}\left(\frac{c_{\mathrm{s}}}{0.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)\,\mathrm{Myr},$ (1)
for the stars with mass $m_{*}$ with disc of scale radius $R_{1}$, at which the Keplerian frequency is $\Omega_{1}$. Equation 1 applies for discs for which accretion from the disc onto the star is driven by turbulent viscosity, while the environmental dependence would enter via $\alpha_{\mathrm{SS}}$ if turbulence were driven by ISM acccretion. We generally expect $R_{1}$ to be smaller than the outer radius $R_{\mathrm{CO}}$ by a factor few, so we can estimate $R_{1}=f_{\mathrm{R}}R_{\mathrm{CO}}\approx f_{\mathrm{R}}250\,(m_{*}/1\,M_{\odot})$ au to yield
$\tau_{\mathrm{acc}}\approx 0.2\,\left(\frac{\alpha_{\mathrm{SS}}}{0.1}\right)^{-1}\left(\frac{f_{\mathrm{R}}}{0.3}\right)^{1/2}\left(\frac{m_{*}}{1\,M_{\odot}}\right)\left(\frac{c_{\mathrm{s}}}{0.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)\,\mathrm{Myr}.$ (2)
For surviving discs, the median $\tau_{\mathrm{acc}}$ in our model evolves from $\sim 0.3$ Myr to $\sim 2$ Myr over the course of $\sim 9$ Myr, as discs with short $\tau_{\mathrm{acc}}$ are preferentially dispersed (Extended Data Figure 4). The median stellar mass in our model is $m_{*}\approx 0.2\,M_{\odot}$, and therefore the early median accretion time-scale requires $\alpha_{\mathrm{SS}}\sim 10^{-2}$, reducing to $\alpha_{\mathrm{SS}}\sim 10^{-3}$ at later times. As shown in Figure 4, this median $\alpha_{\mathrm{SS}}$ is consistent with what we infer for the ISM driven turbulence from our model. This range of typical $\alpha_{\mathrm{SS}}$ is also broadly consistent with observational constraints for stellar accretion rates.[7] Within the BHL accretion scenario, we also expect some enhancement of the stellar accretion rate due to capture of material with low angular momentum, or angular momentum misaligned with the existing disc.[66] Overall, we conclude that capture of the ISM is a plausible driver of observed stellar accretion. However, we emphasise that this depends on the unknown $\epsilon_{\mathrm{t}}$, and our finding certainly does not preclude magnetically driven disc winds also contributing to angular momentum transport,[67] particularly in discs with low levels of turbulence.
In summary, we have shown accretion of the ISM is an important process in driving protoplanetary disc evolution. In particular, our three most important conclusions are:
  1. 1.
    Disc masses, radii and stellar accretion rates as a function of stellar mass and age can result directly from Bondi-Hoyle-Lyttleton accretion.
  2. 2.
  3. 3.
0.1 Acknowledgements:
We thank Cathie Clarke for useful discussion. AJW has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101104656. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (PROTOPLANETS, grant agreement No. 101002188).
0.2 Competing interests:
The authors declare no competing interests.
0.3 Materials & Correspondence:
Requests for materials or correspondence should be addressed to A.J.W., currently at andrew.winter@oca.eu.
Figure 1: Local velocity, density and normalised Bondi-Hoyle-Lyttleton accretion rates. The relative velocity (top) and density (middle) of the local gas, for a subset of six stars in our sample of 500 trajectories. The BHL accretion rate (bottom) is normalised to be independent of mass $m_{*}$, by which lines are coloured. The gas density and velocity are shown at the accretion scale (see Methods). In the middle panel, the horizontal dotted lines are the mean gas density for some high density regions in Taurus (see Methods).[68] In the bottom panel, the red solid (dashed) line follows the median (mean) normalised BHL accretion rate for the entire sample, and a normalised two-dimensional histogram.
Figure 2: Disc properties resulting from ISM accretion compared to observations. Linear kernel density estimates (KDEs) for the disc mass (top), stellar accretion rate (middle) and accretion radii (bottom) versus stellar mass in our model compared to observations (red points, red dashed lines) in local star forming regions. We show results at times $0.5$, $1$, $2$ and $3$ and $5$ Myr. The red points represent the distribution of 100×100$\times$100 × dust disc masses (top), stellar accretion rate (middle) in $\rho$ Ophiuchus, Taurus, Lupus, Chameleon I and Upper Scorpius (left to right).[42] When constructing the KDEs, model points are weighted to yield the same distribution along the stellar mass axis as in the observational samples, except in the bottom row. For comparison with the accretion radii, we show the empirically inferred relationship between CO outer disc radii as a red dashed line in the bottom panels.[48]
Figure 3: Half-life replenishment factor evolution. The fraction of mass that has been accreted in the last half of the disc life-time, $f_{M,1/2}$, as a function of time for $500$ discs. On the left, we show the time evolution of $f_{M,1/2}$ for a random subset of $100$ discs for clarity. These lines are coloured by the ratio of the disc mass to stellar mass $M_{\mathrm{disc}}/m_{*}$, shown in the colour bar along the top. On the right hand side we show a cumulative fraction for $f_{M,1/2}$ at $1$, $2$, $3$ and $8$ Myr. Only discs with a total mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ are included in this cumulative fraction, lower mass discs are assumed to be dispersed.
Figure 4: Turbulent viscosity driven by in-fall. Evolution of the cumulative distribution function for the turbulent Mach number (or viscous $\alpha_{\mathrm{SS}}$, top axis) for the discs in our model, assuming a maximal energy input efficiency $\epsilon_{\mathrm{t}}=1$. Each line colour represents a different disc age in our model, from $0.1$ to $9$ Myr. We show as coloured vertical lines some examples of discs with constraints on their turbulent velocity from the literature, several of which are upper limits.[7]
METHODS
0.4 Overview:
@@
Our aim is to estimate the distribution of accretion histories for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. To do so, we must follow the density and relative velocity of gas surrounding these young stars. This includes both the period in which it is bound to the gaseous overdensity in which it forms,[36] and once the bound gas has dispersed. The most direct approach would be to perform full hydrodynamic simulations of star formation in the turbulent interstellar medium within the galaxy. This would require capturing star formation physics over a wide range of scales, maintaining the turbulent cascade that propagates from the galactic[37, 69, 70] down to the protoplanetary disc size scale for the several Myr over which protoplanetary discs persist. This is computationally challenging, if not impracticable. However, we can still estimate the statistical evolution of the accretion rate semi-analytically by applying an excursion set formalism.[39] This method is effectively zero dimensional, and is therefore a highly efficient way of statistically following density fluctuations of the ISM.
@@
    Stellar dynamics in the bound cloud: Draw a representative stellar population, using the star formation rate based on step 2 to statistically calculate when stars should form during the collapse of a region. Compute the dynamical evolutioon of these stars under the cloud potential until cloud dispersal, tracking the local gas density and velocity.
@@
    Accretion rate onto the star-disc system: Calculate the Bondi-Hoyle-Lyttleton (BHL) accretion rate onto the star-disc system over time.
@@
In the remainder of this work, we describe these steps in greater detail. We then delve deeper into some of the implications of our model.
0.5 Properties of local star forming regions (step 1):

### v2

@@ Abstract
Planet formation occurs over a few Myr within protoplanetary discs of dust and gas, which are often assumed to evolve in isolation. However, extended gaseous structures have been uncovered around many protoplanetary discs, suggestive of late-stage in-fall from the interstellar medium (ISM). To quantify the prevalence of late-stage in-fall, we apply an excursion set formalism to track the local density and relative velocity of the ISM over the disc lifetime. We then combine the theoretical Bondi-Hoyle-Lyttleton (BHL) accretion rate with a simple disc evolution model, anchoring stellar accretion time-scales to observational constraints. Disc lifetimes, as well as masses, stellar accretion rates and approximate gas radii as a function of stellar mass and age are remarkably well-reproduced by our simple model that includes only ISM accretion. If BHL accretion is the origin of the initial disc material, then $20{-}70$ percent of discs are mostly composed of material accreted in the most recent half of their lifetime, suggesting that disc properties are not a direct test of isolated evolution models. The ISM also deposits sufficient energy to drive observed turbulence in protoplanetary discs with viscous $\alpha_{\mathrm{SS}}\sim 10^{-5}-10^{-1}$, although we emphasise that angular momentum transport may still be driven internal processes. Overall, our findings represent a compelling case for regulation of planet formation by galactic-scale turbulence, with broad consequences for planet formation theory. Our simple approach can be easily applied to semi-analytic models, and our results represent urgent motivation for an unbiased survey to confirm or refute our findings.
protoplanetary discs – planet formation – star forming regions
1 Introduction
Planets form and begin their evolution within discs of gas and dust that persist for a few Myr around young stars (e.g. Haisch et al., 2001). The properties of these protoplanetary discs, and the physics that shapes them, is critical to our understanding of the origin of the (exo)planets we observe today. For many years, models for planet formation have been based on assumptions that the star-disc system by-and-large evolves in isolation from its surroundings. This assumption is now being challenged on a variety of fronts. External mechanisms that shape protoplanetary discs include irradiation from nearby massive stars (Winter & Haworth, 2022, and references therein), neighbour star-disc encounters (Cuello et al., 2023, and references therein) and the late stage in-fall of gas from the surroundings (e.g. Dullemond et al., 2019; Kuffmeier et al., 2020, 2021, 2023). It is this last influence that is the topic of this work.
@@ 1 Introduction
Recent observations in molecular line emission (Huang et al., 2021) and infrared scattered light (Garufi et al., 2020; Benisty et al., 2023) are frequently revealing extended structures associated with protoplanetary discs, on spatial scales several times the disc outer radius. In-fall has been suggested as a possible origin for these structures (Dullemond et al., 2019; Kuffmeier et al., 2020, 2023; Hanawa et al., 2024), which are highly complex, including large scale spirals and streamers around T Tauri (Huang et al., 2023) and Herbig AeBe (Boccaletti et al., 2020) stars, often associated with reflection nebulosity (Gupta et al., 2023), accretion outbursts (Pérez et al., 2020; Hales et al., 2020; Dong et al., 2022) and misaligned inner discs (Ginski et al., 2021). Evidence that is suggestive of in-fall events has been found not only during the earliest stages of disc evolution, when the star is still embedded in its natal environment (Jørgensen et al., 2009; Maury et al., 2019; Codella et al., 2024), but also around stars such as AB Aur (Nakajima & Golimowski, 1995; Grady et al., 1999; Fukagawa et al., 2004) DR Tau (Mesa et al., 2022) and SU Aur (Ginski et al., 2021), which are all mature star-disc systems of age $\sim 1{-}3$ Myr. A recent estimate from a sample size of $43$ disc-hosting stars suggests approximately $16$ percent are interacting with the ambient ISM (Garufi et al., 2024). This material should fall onto the disc on a time-scale much shorter than the disc lifetime (Dullemond et al., 2019; Kuffmeier et al., 2020), suggesting that either these events occur much more frequently than observed or the external structure is replenished by the ISM for a prolonged period.
In-falling gas may be the consequence of Bondi-Hoyle-Lyttleton (BHL) accretion, where low velocity gas is captured due to the gravitational potential of the star (Bondi, 1952; Edgar, 2004). Theoretically, the rate of capture of material from the interstellar medium (ISM) via BHL accretion can be substantial while stars remain bound to a dense gas reservoir. Padoan et al. (2005) proposed this mechanism to explain the observed relationship between stellar accretion rates and stellar mass. This idea was supported by subsequent work (Throop & Bally, 2008; Klessen & Hennebelle, 2010), while more recent analysis of turbulent hydrodynamic simulations in a periodic box have shown that late stage accretion may be significant source of mass and angular momentum for $\sim 1$ Myr old discs (Kuffmeier et al., 2023; Pelkonen et al., 2024; Padoan et al., 2024). However, assessing the importance of this process for the observed population of protoplanetary discs throughout their lifetime requires quantifying the temporal evolution of the turbulent ISM surrounding forming stars. This is challenging for hydrodynamic simulations that are both expensive and temporally limited by the turbulent crossing time of the computational domain.
@@
In this work, we present a semi-analytic approach for tracking the BHL accretion rate over the disc lifetime. In a sense, this is a complement and extension of that of Throop & Bally (2008), who tracked BHL accretion rates onto stars that are gravitationally bound to a dense gas reservoir. Our aim here is to estimate the distribution of accretion histories over the whole disc lifetime for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. This includes both the period in which the star is bound to the gaseous overdensity in which it forms (Throop & Bally, 2008) and once the bound gas has dispersed. Statistically interpreting the role of BHL accretion on a disc population level requires simulating the turbulent cascade from galactic to protoplanetary disc size scales for the several Myr over which discs persist (Klessen & Hennebelle, 2010; Colman et al., 2022; Nusser & Silk, 2022). To do so, we apply an excursion set formalism (Hopkins, 2012) which is effectively zero dimensional. This is therefore a highly efficient way of statistically following density fluctuations of the ISM, making it conducive to semi-analytic planet formation models (e.g. Weder et al., 2023; Mordasini & Burn, 2024).
@@
    Stellar dynamics in the bound cloud: Draw a representative stellar population, using the star formation rate based on step 2 to statistically calculate when stars should form during the collapse of a region. Compute the dynamical evolution of these stars under the cloud potential until cloud dispersal, tracking the local gas density and velocity.
@@
    Accretion rate onto the star-disc system: Calculate the BHL accretion rate onto the star-disc system over time.
@@
In the remainder of this work, we describe these steps in greater detail, and discuss the implications for protoplanetary disc evolution.
2 Modelling the Interstellar Medium
2.1 Overview
In this section, our aim is to estimate the distribution of accretion histories for a population of young stars that resembles those surveyed for disc properties in the solar neighbourhood. To do so, we must follow the velocity and density distribution of gas surrounding these young stars. This includes both the period in which it is bound to the gaseous overdensity in which it forms (cf. Throop & Bally, 2008), and once the bound gas has dispersed. We estimate the statistical evolution of the accretion rate semi-analytically by applying an excursion set formalism (Hopkins, 2012). This method is effectively ‘0D’, and is therefore a highly efficient way of statistically following density fluctuations of the ISM.
We adopt the following approach in the remainder of this section. First, we introduce the excursion set formalism and apply it to review the typical properties of local giant molecular clouds (GMCs) and star forming regions (Section 2.2). We then introduce a simple model for star formation during the collapse of bound GMCs (Section 2.3), and the approach for modelling the dynamics of stars forming in this environment (Section 2.4). We consider the temporal evolution of density and velocity of the unbound, galactic ISM outside of the star forming region in Section 2.5. Finally, we evaluate how the accretion rate for individual stars evolves as they form from the local population of GMCs (Section 2.6).
2.2 Properties of local star forming regions
Figure 1: Sampling of the mass function via the excursion set formalism for GMCs (blue points) based on $10^{5}$ randomly drawn density perturbation trajectories in the solar neighbourhood. We have normalised the $y$-axis such that we show the average number of regions younger than $10$ Myr that are expected to be found within $200$ pc. We highlight in yellow the region which is well-sampled, corresponding to typical local star forming regions which are surveyed for protoplanetary disc properties. In cyan, we show the range of initial GMC masses we consider to generate our disc population synthesis. In orange, we show the same mass function assuming a fixed star formation efficiency $\epsilon=0.5$.